# Multi-contrast laser endoscopy for in vivo gastrointestinal imaging

### v1

@@ Results
Diffuse reflectance spectroscopy has been widely studied as a non-contact, label-free method for revealing differences in underlying tissue makeup and structure. To evaluate the accuracy of multispectral reflectance measurements acquired with MLE, we imaged a color calibration chart with 18 patches that exhibit distinct spectral reflectance signatures. Compared to clinical WLE and NBI, which capture two to three bands of spectral information (Fig. 2a), MLE measures spectral reflectance at 8 distinct wavelengths (Fig. 2b). Qualitative inspection of the spectral reflectance values measured by MLE and a reference spectrometer shows strong agreement for each Macbeth patch (Fig. 2c). Quantitative comparison with the reference spectra results in a mean absolute reflectance error of 0.05 $\pm$ 0.01, with each spectral channel error centered near 0 (Fig. 2d).
@@
In addition to the low-coherence laser diodes (MLE-LC) used for spectral imaging, a high-coherence laser (MLE-HC) was integrated into the multi-contrast illumination source to produce well-defined speckle for laser speckle contrast imaging (LSCI). As shown in Fig 3a, MLE-LC illumination produce images with speckle contrast ($\mathrm{K}$) similar to clinical WLE, while MLE-HC illumination produces images with an order of magnitude greater speckle contrast.
@@
To assess the sensitivity of MLE to variations in blood flow in the presence of motion in vivo, we imaged arterioles and capillaries in the human soft palate mucosa in-vivo (Fig 3f). The flow contrast was improved by registering sequential speckle contrast frames, via interlaced color imaging, and averaging them (see Methods). Qualitative inspection of the flow contrast maps shows that the vasculature becomes distinguishable from the background tissue with an average of 10 or more frames (Fig 3g). To quantify this improvement, the root-mean-square (RMS) contrast [26] between the vasculature and background regions of interest (ROI) was calculated for each temporal window size (Fig 3h). These results demonstrate that RMS contrast increases significantly over the first 0-7 frames, then gradually improves as the number of frames increases.
@@
We also explored using functional measurements as a source of image contrast. The reflectance images from MLE were used to spatially map total oxygen saturation (Fig. 5e). The average total oxygen saturation ($\overline{x}$=51.3%) is consistent with prior work [29], but the differences between polyp and normal tissue were not statistically significant (Fig. 5j). Laser speckle illuminated frames (Fig. 5f) were acquired to spatially map blood flow with LSCI (Fig. 5g). The mean blood flow measured with LSCI was 9.0% less in polyp tissue than in normal tissue, and the difference in average blood flow for polyp and normal regions was found to be statistically significant (Fig.5k). In some samples, the decrease in blood flow was relatively uniform within the margins of the polyps, making them visually identifiable, whereas in other cases, the blood flow was non-uniform in appearance. The poor distinction between tissue types is reflected in the quantitative analysis, where the modalities have similar RMS contrast values as WLE and NBI (Fig.5n).
@@ Methods
Color images were simulated using the spectral reflectance measurements, the transmission spectrum of a Bayer camera filter [44], and the measured output illumination spectra from the clinical light source. Optimized weights for the simulated spectrally enhanced images were identified using gradient descent optimization, with an objective function designed to maximize the CIEDE2000 color difference ($E\delta_{00}$[45] between 100 pixels sampled from normal and polyp tissue regions in each image. WLE and NBI images were also simulated, allowing the CIEDE2000 color difference to be computed for the same pixels as the spectral enhanced images. Prior to processing, the reflectance channels for each wavelength were rescaled to the average intensity for that wavelength across the dataset. Each pixel was normalized so that its maximum color channel value equaled 0.8 to remove the effect of brightness in the color difference calculation [46]. For visualization, the images were scaled so that the mean intensity of the image was equal to 0.4. See Supplementary Fig. 7) for additional implementation details.
@@ Acknowledgments
This work was supported in part with funding from the National Institutes of Health Trailblazer Award (R21 EB024700), the National Science Foundation (NSF) Graduate Research Fellowship Program (DGE-1746891), and the NSF CAREER Award No. 2146333. The authors thank Terry Mistalski (HCS Solutions Pro), Christopher Perry (Cantel Medical Corporation), Tom Arcand (Cantel Medical Corporation), Daniel Pferdeort (Cantel Medical Corporation), Kiona Williams (Johns Hopkins Hospital Central Sterile Processing), and Elaine Nowakowski (Johns Hopkins Hospital Epidemiology and Infection Control) for assistance with infection prevention measures. The authors also thank Greg McKay for sharing photographs of the system for inclusion in Fig. 1.

### v2

@@ Results
Diffuse reflectance spectroscopy has been widely studied as a non-contact, label-free method for revealing differences in underlying tissue makeup and structure. To evaluate the accuracy of multispectral reflectance measurements acquired with MLE, we imaged a color calibration chart with 18 patches that exhibit distinct spectral reflectance signatures. Compared to clinical WLE and NBI, which capture two to three bands of spectral information (Fig. 2a), MLE measures spectral reflectance at 8 distinct wavelengths (Fig. 2b). Qualitative inspection of the spectral reflectance values measured by MLE and a reference spectrometer shows strong agreement for each Macbeth patch (Fig. 2c). Quantitative comparison with the reference spectra results in a mean absolute reflectance error of 0.04 $\pm$ 0.03, with each spectral channel error centered near 0 (Fig. 2d).
@@
In addition to the low-coherence laser diodes (MLE-LC) used for spectral imaging, a high-coherence laser (MLE-HC) was integrated into the multi-contrast illumination source to produce well-defined speckle for laser speckle contrast imaging (LSCI). As shown in Fig. 3a, MLE-LC illumination produce images with speckle contrast ($\mathrm{K}$) similar to clinical WLE, while MLE-HC illumination produces images with an order of magnitude greater speckle contrast.
@@
To assess the sensitivity of MLE to variations in blood flow in the presence of motion in vivo, we imaged arterioles and capillaries in the human soft palate mucosa in-vivo (Fig. 3f). The flow contrast was improved by registering sequential speckle contrast frames, via interlaced color imaging, and averaging them (see Methods). Qualitative inspection of the flow contrast maps shows that the vasculature becomes distinguishable from the background tissue with an average of 10 or more frames (Fig. 3g). To quantify this improvement, the root-mean-square (RMS) contrast [26] between the vasculature and background regions of interest (ROI) was calculated for each temporal window size (Fig. 3h). These results demonstrate that RMS contrast increases significantly over the first 1-6 frames, then gradually improves as the number of frames increases.
@@
We also explored using functional measurements as a source of image contrast. The reflectance images from MLE were used to spatially map total oxygen saturation (Fig. 5e). The average total oxygen saturation ($\overline{x}$=50.1$\pm$ 6.9%) is consistent with prior work [29], but the differences between polyp and normal tissue were not statistically significant (Fig. 5j). Laser speckle illuminated frames (Fig. 5f) were acquired to spatially map blood flow with LSCI (Fig. 5g). The mean blood flow measured with LSCI was less in polyp tissue than in normal tissue ($\overline{x}$=-10.0$\pm$ 20.7%), but the difference was not statistically significant (Fig. 5k). In some samples, the decrease in blood flow was relatively uniform within the margins of the polyps, making them visually identifiable, whereas in other cases, the blood flow was non-uniform in appearance. The poor distinction between tissue types is reflected in the quantitative contrast analysis, where the modalities have similar RMS contrast values as WLE and NBI (Fig. 5m).
@@ Methods
Color images were simulated using the spectral reflectance measurements, the transmission spectrum of a Bayer camera filter [44], and the measured output illumination spectra from the clinical light source. Optimized weights for the simulated spectrally enhanced images were identified using gradient descent optimization, with an objective function designed to maximize the CIEDE2000 color difference ($E\delta_{00}$[45] between 100 pixels sampled from normal and polyp tissue regions in each image. WLE and NBI images were also simulated, allowing the CIEDE2000 color difference to be computed for the same pixels as the spectral enhanced images. Prior to processing, the reflectance channels for each wavelength were rescaled to the average intensity for that wavelength across the dataset. Each pixel was normalized so that its maximum color channel value equaled 0.8 to remove the effect of brightness in the color difference calculation [46]. For visualization, the images were scaled so that the mean intensity of the image was equal to 0.4. See Supplementary Fig. 7 for additional implementation details.
@@ Acknowledgments
This work was supported in part with funding from the National Institutes of Health Trailblazer Award (R21 EB024700), the National Science Foundation (NSF) Graduate Research Fellowship Program (DGE-1746891), and the NSF CAREER Award No. 2146333. The authors thank Terry Mistalski (HCS Solutions Pro), Christopher Perry (Cantel Medical Corporation), Tom Arcand (Cantel Medical Corporation), Daniel Pferdeort (Cantel Medical Corporation), Kiona Williams (Johns Hopkins Hospital Central Sterile Processing), and Elaine Nowakowski (Johns Hopkins Hospital Epidemiology and Infection Control) for assistance with infection prevention measures. The authors also thank Greg McKay for sharing photographs of the system included in Fig. 1.